# Ridge Regression for Paired Comparisons: A Tractable New Approach, with Application to Premier League Football

### v1

@@ Abstract
Paired comparison models, such as Bradley-Terry and Thurstone-Mosteller, are commonly used to estimate relative strengths of pairwise compared items in tournament-style datasets. With predictive performance as primary criterion, we discuss estimation of paired comparison models with a ridge penalty. A new approach is derived which combines empirical Bayes and composite likelihoods without any need to re-fit the model, as a convenient alternative to cross-validation of the ridge tuning parameter. Simulation studies, together with application to 28 seasons of English Premier League football, demonstrate much better predictive accuracy of the new approach relative to ordinary maximum likelihood. While the application of a standard bias-reducing penalty was found to improve appreciably the performance of maximum likelihood, the ridge penalty with tuning as developed here yields greater accuracy still.
@@ 1 Introduction
The problem of rating a list of items on the basis of a set of paired comparisons arises frequently, in a variety of fields including artificial intelligence (e.g., Chen et al., 2013; Rafailov et al., 2023), bibliometrics (Stigler, 1994; Varin, Cattelan and Firth, 2016), education (e.g., Wheadon et al., 2020; Bartholomew and Jones, 2022), forensic science (Thompson et al., 2018), genetics (e.g., Ma, Wong and Owen, 2012), politics (e.g., Loewen, Rubenson and Spirling, 2012), psychometrics (e.g., Maydeu-Olivares and Böckenholt, 2005) and sport (e.g., Glickman and Stern, 2016), to name just a few. A classical reference for the statistical analysis of paired comparisons is David (1988). A more recent review is Cattelan (2012), while Aldous (2017) discusses various aspects of paired comparisons from the point of view of applied probability.
Although the application areas amenable to paired-comparison modelling are many, in what follows we adopt the language of sport; this is due to the football application that motivated this work and which is described in Section 5. Consider $p$ teams. In paired comparison models, the probability that team $i$ beats team $j$ in a match between the two teams is
@@
where $F(\cdot)$ is a the cumulative distribution function of a zero-symmetric continuous random variable and $\mu_{i}$ is the strength of team $i$. Popular choices for $F$ are the logistic distribution which gives the Bradley-Terry model (Bradley and Terry, 1952) and the normal distribution which corresponds to the Thurstone-Mosteller model (Thurstone, 1927; Mosteller, 1951). These two models are quite similar in practice, given the well-known correspondence between the logit link assumed in the Bradley-Terry model and the probit link of the Thurstone-Mosteller model (e.g., Agresti, 2002, pages 246–247).
We begin by assuming that the outcome of a match is independent of the order in which the teams are presented, in other words that there is no home-field effect; and also that there are no ties. These assumptions will be relaxed later in Section 3.2.
Maximum likelihood estimation of the vector of strength parameters $\bm{\mu}=(\mu_{1},\ldots,\mu_{p})^{\top}$ is well known to be problematic, in part because the estimate of $\mu_{i}$ diverges if team $i$ wins or loses all of its matches (e.g., Kosmidis and Firth, 2021). Bias-reduced maximum likelihood (Firth, 1993) overcomes the limitations of maximum likelihood, guaranteeing finiteness and also better frequentist properties of the strength estimates (Kosmidis and Firth, 2021). Finite estimates can also be obtained through other forms of penalized maximum likelihood (Mease, 2003) or Bayesian methods (Caron and Doucet, 2012).
@@ 2 Ridge Regression for Paired Comparisons
Denote the outcome of the match between teams $i$ and $j$ as $Y_{ij}=1$ if $i$ wins and $Y_{ij}=-1$ if $j$ wins. In this paper we investigate a natural form of regularization for estimation of the strength parameters, based on the ridge-penalized log likelihood
@@
Here $\mathcal{S}$ denotes the played tournament, i.e., $\mathcal{S}$ is the set of index pairs $(i,j)$ in which $y_{ij}$ is observed and $\lambda\geq 0$ a scalar tuning parameter. For a fixed value of $\lambda$, computation of the ridge estimates of the strength parameters is a relatively simple optimization that can be carried out, for example, with standard quasi-Newton algorithms. The standard approach to determine a good value for $\lambda$ is cross-validation. The choice of the splits used in cross-validation requires care in the case of incomplete tournaments to guarantee that each assessment set is comparable to the source data. This work is motivated by the estimation of team strengths during ongoing round-robin tournaments, and in this case appropriate balance in the assessment sets is obtained by using match-weeks as cross-validation units.
An alternative to cross-validation is estimation of $\lambda$ using the empirical Bayes method (Morris, 1983). Ridge regression coincides with maximum-a-posteriori estimation under the assumption that the strength parameters are uncorrelated zero-mean normal random variables with precision $\lambda$,
@@
where ${\text{I}}_{p}$ is the identity matrix of size $p$. The empirical Bayes recipe estimates $\lambda$ by maximizing the marginal likelihood obtained by integrating out the team strengths. This leads to a cumbersome integral of dimension $p$,
@@
3 The Pairwise Empirical Bayes Method
@@ 3.1 Estimation of the tuning parameter
We develop here a simple method that avoids the potentially high-dimensional integration (5) by using composite likelihood methods (Lindsay, 1988; Varin, Reid and Firth, 2011). The tuning parameter $\lambda$ is estimated by maximizing the pairwise log-likelihood constructed from all pairs of correlated matches, i.e., those pairs with one team in common, such as ($i$ vs $j$) and ($i$ vs $k$). The pairwise log-likelihood for $\lambda$ is
$\ell_{\text{pair}}(\lambda)=\sum_{r,s\in\mathcal{O}}n_{rs}\log p_{rs}(\lambda),$ (6)
where $\mathcal{O}=\{-1,+1\}$ is the set of possible match outcomes (which will be later extended in Section 3.2 to include ties), $n_{rs}=\sum_{ijk}1(Y_{ij}=r,Y_{ik}=s)$ is the number of correlated pairs of matches with outcomes $r$ and $s$ and $p_{rs}(\lambda)=\mathrm{Pr}\left(Y_{ij}=r,Y_{ik}=s\right)$ is the corresponding marginal bivariate probability.
In the following we use the Thurstone-Mosteller model, because it is mathematically more convenient than the Bradley-Terry model. Under the Thurstone-Mosteller model, the bivariate probabilities $p_{rs}$ have closed-form expressions by a well-known result sometimes referred to as the Theorem of Median Dichotomy (Sheppard, 1898; Cox and Wermuth, 2002). The probability of two successes for the common player $i$ is
$p_{11}(\lambda)=\int_{0}^{\infty}\int_{0}^{\infty}\phi_{2}\left(u,v;\frac{\lambda^{-1}}{1+2\lambda^{-1}}\right)\text{d}u\text{d}v=\frac{(1+\tau)}{4},$
@@
takes a value in the interval $(0,1/3)$. The general expression for $p_{rs}$ depends on whether there is concordance or not in the outcomes of the two matches:
@@
$\ell_{\text{pair}}(\lambda)=c\log(1+\tau)+d\log(1-\tau),$
where $c$ is the count of concordant correlated pairs (win-win, loss-loss) and $d$ the count of discordant correlated pairs (win-loss, loss-win). The maximum pairwise likelihood estimator of $\tau$ is then the familiar Kendall $\tau$ coefficient
@@
In sparse tournaments, where the number of matches per team is small, it is convenient to consider the classical adjustment to contingency tables that consists here in adding one imaginary concordant pair and one imaginary discordant pair to the counts for each of the $p$ teams. The corresponding small-sample adjusted estimator of $\tau$ is then
$\hat{\tau}=\frac{c-d}{c+d+p},$
which coincides with the maximizer of the penalized pairwise log-likelihood
~$\tilde{\ell}_{\text{pair}}(\lambda)={\ell}_{\text{pair}}(\lambda)+p\log(1-\tau^{2}).$ (9)
3.2 Home-field effect and ties
The order of presentation of the teams or other items is important in many paired comparison contexts, notable examples being the home-field effect in sport competitions or the white player advantage in chess. In paired-comparison models, such an order effect is typically described through the inclusion of an intercept term (e.g., Agresti, 2002),
$\Pr(Y_{ij}=1|\bm{\mu})=\Phi(\delta+\mu_{i}-\mu_{j}),$ (10)
@@ 3.2 Home-field effect and ties
Another common complication in paired comparisons is the presence of ties, represented by the null outcome $Y_{ij}=0$ whenever $i$ and $j$ compete but neither of them wins. As discussed in Agresti (1992), ties can be incorporated in paired comparison models through a cumulative link model. The Thurstone-Mosteller model with home-field effect corresponds to the latent continuous process $Z_{ij}=\delta+\mu_{i}-\mu_{j}+\epsilon_{ij},$ where $\epsilon_{ij}$ are uncorrelated standard normal variables. The outcome of the comparison depends on a threshold parameter $\gamma\geq 0$, so that $j$ beats $i$ if $Z_{ij}<-\gamma$, $i$ ties with $j$ if $-\gamma\leq Z_{ij}<\gamma$ and $i$ beats $j$ if $Z_{ij}\geq\gamma.$ The corresponding cumulative probabilities for the three outcomes given the strengths are
$\Pr(Y_{ij}\leq x|\bm{\mu})=\Phi(c_{x}-\mu_{i}+\mu_{j}),\quad x\in\{-1,0,1\},$
@@
As before, we proceed by empirical Bayes estimation of the model parameters $\gamma$, $\delta$ and $\lambda$ from the marginal distribution of the matches, and then compute the ridge estimates of the strength parameters.
The marginal probabilities of losing, drawing and winning for the home team are simply calculated by setting all strength parameters to zero, giving $\Pr(Y_{ij}=x)=\Phi(c_{x})-\Phi(c_{x-1})$, for $x\in\{-1,0,1\}$. From these marginal probabilities, consistent estimates of $\gamma$ and $\delta$ are readily obtained as
$\hat{\gamma}=\frac{1}{2}\left\{\Phi^{-1}(1-\hat{p}_{1})-\Phi^{-1}(\hat{p}_{-1})\right\},\quad\hat{\delta}=\frac{1}{2}\left\{\Phi^{-1}(\hat{p}_{1})-\Phi^{-1}(\hat{p}_{-1})\right\},$ (11)
where $\hat{p}_{-1}=\sum_{ij}1(Y_{ij}=-1)/(n+1)$ and $\hat{p}_{1}=\sum_{ij}1(Y_{ij}=1)/(n+1)$ are the sample estimates of the probability of a win away and at home, respectively. In tournaments without ties, $\hat{\delta}=0$ as expected.
Having estimated $\delta$ and $\gamma$, we proceed with estimation of the tuning parameter $\lambda$ by maximizing the pairwise log-likelihood (6) where $p_{rs}(\lambda)$ now denotes the joint probability of the outcomes of two matches with either the same home team or the same away team. For example, the probability of a victory and a tie in two matches played at home by $i$ is
@@
The other bivariate probabilities involved in the pairwise likelihood are similarly computed. In these probabilities $\gamma$ and $\delta$ are replaced with their estimates (11) so that the pairwise likelihood depends on $\lambda$ only.
Maximization of the pairwise likelihood for $\lambda$ is conveniently performed by using the robust algorithm of Brent (1973, ch. 4) on the finite interval domain of $\tau$ from equation (7), with subsequent back-transformation as in equation (8). The same small-sample adjustment as before remains available through the pairwise likelihood penalty of equation (9).
@@ 4.1 True strengths normally distributed
The performance of the pairwise empirical Bayes method is illustrated by simulations from the Thurstone-Mosteller model with normally distributed strength parameters. We considered 120 scenarios given by the combination of: (1) three values for the number of teams $p\in\{20,40,60\}$, (2) eight values for the tuning parameter $\lambda\in\{2^{-1},2^{0},\ldots,\allowbreak 2^{6}\}$ and (3) five sample sizes corresponding to five increasing fractions of the matches in a double round-robin tournament. More specifically, we simulated the full double-round-robin tournament and used the first $m$ match-weeks for training and the remaining $2(p-1)-m$ match-weeks for testing. The number of training weeks was chosen so as to approximate a specified proportion of the tournament matches. For example, when we consider 20% of matches for training and the number of teams is $p=30$, then the training set is taken to be the first 12 match-weeks because $0.2\times(2\times 29)=11.6$. The five sample fractions considered for training in our simulations are $20\%,30\%,40\%,50\%$ and $80\%$ of the matches in the double round-robin tournament. For each scenario, $1000$ replications were generated and analysed.
@@
The performance of our pairwise empirical Bayes method is compared here with: (1) maximum likelihood estimation, (2) bias-reduced maximum likelihood and (3) ridge estimation with $\lambda$ estimated by cross-validation. Bias-reduced maximum likelihood estimation is computed through the R (R Core Team, 2024) package brglm2 (Kosmidis, 2021). Since paired-comparison models are specified through pairwise differences of strengths, an arbitrary constraint is needed for identification of the maximum likelihood and bias-reduced maximum likelihood estimates. In the simulations we used the constraint $\mu_{1}=0$. Cross-validated ridge estimates were calculated via the glmnet package (Friedman, Tibshirani and Hastie, 2010; Tay, Narasimhan & Hastie, 2023), using match-weeks as the assessment unit. Paraphrasing the typical language used in model validation, we could call such a design a ‘leave one match-week out’ cross-validation.
@@
We simulated matches without ties because we decided to focus on the estimation of the strength parameters. Estimation of the tie threshold parameter $\gamma$ would not be problematic, but bias-reduced maximum likelihood and cross-validated logistic regression are easier to compute in paired comparison models without ties. In all simulations the home-field effect parameter was set to $\delta=0.2$, so that approximately 58% of matches are won by the home team since $\Phi(0.2)\approx 0.58$.
@@
The predictive performance of each method is measured by the classical logarithmic score LS (Good, 1952), defined as the negative average log-likelihood for the testing matches computed at the strengths estimated with the training matches. As reference we consider the naive forecast that predicts the result of every match using the assumed probability of a home win, that is $\text{LS}_{\text{naive}}=-0.58\log(0.58)-0.42\log(0.42)=0.68$. The naive forecast corresponds to predictions for the paired comparison model with strength parameters all zero, $\mu_{i}=0$ for all $i=1,\ldots,p$. With this reference, we define the logarithmic skill score as
$\text{LSS}=1-\frac{\text{LS}}{\text{LS}_{\text{naive}}}=1-\frac{\text{LS}}{0.68}.$
The higher the logarithmic skill score, the better the predictive performance. Better forecasts on average than those obtained by the naive forecast have a logarithmic skill score above zero.
@@
Figure 1: Logarithmic skill scores computed for incomplete double round-robin tournaments of different sizes. The columns correspond to various levels of incompleteness (20%, 30%, 50% or 80% of all possible matches used for training and the rest of matches for prediction). The rows correspond to different numbers of teams (20, 40, 60). The team strengths are drawn from a zero-mean normal distribution with precision $\lambda$. The strengths are estimated by bias reduced maximum likelihood (BRMLE), maximum likelihood (MLE) and pairwise empirical Bayes (PEB). The reference method is a naive forecast that uses only the expected probability of win and corresponds to a zero logarithmic skill score (grey solid line).
We now comment on the results reported in Figure 1. As expected, as the number of teams or the fraction of matches used for training increases, all methods converge to the same predictive performance. Furthermore, as $\lambda$ increases, estimates of team strengths converge to zero and therefore all methods produce predictions that are on average equivalent to those of the naive forecast. The predictive performance of maximum likelihood estimation is completely unsatisfactory when the training set is small and the number of teams is 20, with its average logarithmic skill scores being worse than the naive forecast. This result is relevant for us because twenty is the number of teams in various major football leagues, such as the English Premier League which motivated this research, which is considered later in Section 5. The bias-reduced maximum likelihood estimator improves substantially on maximum likelihood in most settings, but with small training sets and large values of $\lambda$ it still leads to worse logarithmic skill scores on average than the naive forecast.
As shown in Figure 1, pairwise empirical Bayes has uniformly the best predictive performance of all methods considered. Pairwise empirical Bayes was even found to perform relatively well in a still more extreme setting, with $p=20$ teams and only 4 matches per team (approximately 10% of the tournament) used for training; that extreme setting is not displayed in Figure 1 because maximum likelihood performs so poorly that its logarithmic skill score values fall too far below zero.

### v2

# Tractable Ridge Regression for Paired Comparisons

@@ Abstract
Paired comparison models, such as Bradley-Terry and Thurstone-Mosteller, are commonly used to estimate relative strengths of pairwise compared items in tournament-style data. We discuss estimation of paired comparison models with a ridge penalty. A new approach is derived which combines empirical Bayes and composite likelihoods without any need to re-fit the model, as a convenient alternative to cross-validation of the ridge tuning parameter. Simulation studies demonstrate much better predictive accuracy of the new approach relative to ordinary maximum likelihood. A widely used alternative, the application of a standard bias-reducing penalty, is also found to improve appreciably the performance of maximum likelihood; but the ridge penalty, with tuning as developed here, yields greater accuracy still. The methodology is illustrated through application to 28 seasons of English Premier League football.
@@ 1 Introduction
The problem of rating a list of items on the basis of a set of paired comparisons arises frequently, in a variety of fields including artificial intelligence (e.g., Rafailov et al., 2023), bibliometrics (e.g., Varin et al., 2016), education (e.g., Bartholomew and Jones, 2022), forensic science (e.g., Thompson et al., 2018), genetics (e.g., Ma et al., 2012), politics (e.g., Loewen et al., 2012), psychometrics (e.g., Maydeu-Olivares and Böckenholt, 2005) and sport (e.g., Glickman and Stern, 2017), to name just a few. A classical reference for the statistical analysis of paired comparisons is David (1988). A more recent review is Cattelan (2012), while Aldous (2017) discusses various aspects of paired comparisons from the point of view of applied probability.
In binary paired comparison models, the probability that item $i$ beats item $j$ in a comparison between them is
@@
where $F(\cdot)$ is the cumulative distribution function of a zero-symmetric continuous random variable and $\mu_{i}$ is the strength of item $i$. Popular choices for $F$ are the logistic distribution which gives the Bradley-Terry model (Bradley and Terry, 1952) and the normal distribution which corresponds to the Thurstone-Mosteller model (Thurstone, 1927; Mosteller, 1951). These two models are quite similar in practice, given the well-known correspondence between the logit link assumed in the Bradley-Terry model and the probit link of the Thurstone-Mosteller model (e.g., Agresti, 2002, pages 246–247).
Consider $p$ items. Maximum likelihood estimation of the vector of strength parameters $\bm{\mu}=(\mu_{1},\ldots,\mu_{p})^{\mathsf{T}}$ is well known to be problematic, in part because the estimate of $\mu_{i}$ diverges if item $i$ wins all of its comparisons or loses them all (e.g., Kosmidis and Firth, 2021). Bias-reduced maximum likelihood (Firth, 1993) overcomes such limitations of maximum likelihood, guaranteeing finiteness and also better frequentist properties of the strength estimates (Kosmidis and Firth, 2021). Finite estimates can also be obtained through other forms of penalized maximum likelihood (Mease, 2003) or Bayesian methods (Caron and Doucet, 2012).
The present paper develops a highly tractable alternative regularization of the likelihood for the Thurstone-Mosteller model, using a ridge penalty that is tuned by using a new ‘pairwise empirical Bayes’ method. The predictive performance of this new approach is studied in simulation experiments and through application to match results from incomplete seasons of English Premier League football.
@@ 2 Ridge Regression for Paired Comparisons
We begin by assuming that the outcome of a paired comparison does not depend on the order in which the items are presented, which in sport applications means that there is no home-field effect; and we assume for now also that there are no ties. These assumptions will be relaxed later, in Section 3.2.
Denote the outcome of the paired comparison between items $i$ and $j$ as $Y_{ij}=1$ if $i$ wins and $Y_{ij}=-1$ if $j$ wins. In this paper we investigate a natural form of regularization for estimation of the strength parameters, based on the ridge-penalized log likelihood
@@
Here $\mathcal{S}$ denotes the observed tournament, i.e., $\mathcal{S}$ is the set of index pairs $(i,j)$ in which $y_{ij}$ is observed; and $\lambda\geq 0$ is a scalar tuning parameter. For a fixed value of $\lambda$, computation of the ridge estimates of the strength parameters is a relatively simple optimization, by using for example standard quasi-Newton algorithms. The standard approach to determining a good value for $\lambda$ is cross-validation. An important point is that in pair-comparison data there is often some balancing structure that should be respected in cross-validatory splitting of the dataset; in round-robin tournaments, for example, it will usually make sense to use whole tournament rounds as cross-validation units. When such structure is absent, though, it can often be difficult to identify a suitable set of splits for cross-validation with paired-comparison data.
An alternative to cross-validation is estimation of $\lambda$ using the empirical Bayes method (Morris, 1983). Ridge regression coincides with maximum-a-posteriori estimation when the strength parameters are uncorrelated zero-mean normal random variables with precision $\lambda$,
@@
where ${\text{I}}_{p}$ is the identity matrix of size $p$. Normality of the strengths is purely a working assumption here, to allow determination of $\lambda$ via the empirical Bayes method. In the simulation studies reported in Section 4.2 we show that this method of determining $\lambda$ is robust in the sense that it works well even when the true strengths have appreciably heavier tails than a normal distribution. (There is no need to consider violations of normality due to skewness, because the paired comparison model is identified by differences $\mu_{i}-\mu_{j}$, which have a symmetric distribution in any case.)
The empirical Bayes method determines $\lambda$ by maximizing the marginal likelihood obtained from integrating out the team strengths. This leads to a cumbersome integral of dimension $p$,
@@
3 Pairwise Empirical Bayes
@@ 3.1 Estimation of the tuning parameter
We develop here a simple method that avoids the potentially high-dimensional integration (5) by using composite likelihood methods (Varin et al., 2011). The tuning parameter $\lambda$ is estimated by maximizing the pairwise log-likelihood constructed from all couples of correlated paired comparisons, i.e., those couples with one item in common, such as ($i$ vs $j$) and ($i$ vs $k$). The pairwise log-likelihood for $\lambda$ takes the form
$\ell_{\text{pair}}(\lambda)=\sum_{r,s\in\mathcal{O}}n_{rs}\log p_{rs}(\lambda).$ (6)
Here $\mathcal{O}=\{-1,+1\}$ is the set of possible outcomes of the paired comparisons (which will be later extended in Section 3.2 to include ties); while $n_{rs}=\sum_{ijk}1(Y_{ij}=r,Y_{ik}=s)$ is the number of correlated couples of paired comparisons with outcomes $r$ and $s$, and $p_{rs}(\lambda)=\mathrm{Pr}\left(Y_{ij}=r,Y_{ik}=s\right)$ is the corresponding marginal bivariate probability.
In the following we use the Thurstone-Mosteller model, because it is mathematically more convenient than the Bradley-Terry model and the two models give essentially equivalent fits as already noted in Section 1. Under the Thurstone-Mosteller model, the bivariate probabilities $p_{rs}(\lambda)$ have closed-form expressions, by a well-known result sometimes referred to as the Theorem of Median Dichotomy (Sheppard, 1898; Cox and Wermuth, 2002). The probability of two wins for the common item $i$ is
$p_{11}(\lambda)$ $=\Pr(Y_{ij}=1,Y_{ik}=1)$
$=\int_{0}^{\infty}\int_{0}^{\infty}\phi_{2}\left(u,v;\frac{\lambda^{-1}}{1+2\lambda^{-1}}\right)\text{d}u\text{d}v$
$=\frac{(1+\tau)}{4},$
@@
takes a value in the interval $(0,1/3)$. The general expression for $p_{rs}$ depends on whether there is concordance or not in the outcomes of the two paired comparisons:
@@
$\ell_{\text{pair}}(\tau)=c\log(1+\tau)+d\log(1-\tau),$
where $c$ is the count of concordant correlated couples of paired comparisons ($Y_{ij}=Y_{ik}$) and $d$ the count of discordant couples ($Y_{ij}\neq Y_{ik}$). The maximum pairwise likelihood estimator of $\tau$ is then the familiar Kendall $\tau$ coefficient
@@
In sparse datasets with only a modest number of comparisons per item, it is helpful to consider making a small-sample adjustment along the lines of the classical ‘rule of succession’ of Laplace (1814). In the present context such an adjustment adds one imaginary concordant couple and one imaginary discordant couple per item, and the corresponding small-sample adjusted estimator of $\tau$ is
$\hat{\tau}=\frac{c-d}{c+d+2p}.$
An alternative interpretation of this adjusted estimator is that it maximizes a penalized version of the pairwise log-likelihood, namely
~$\tilde{\ell}_{\text{pair}}(\tau)={\ell}_{\text{pair}}(\tau)+p\log(1-\tau^{2}).$ (9)
3.2 Order effect and ties
The order of presentation of the items is important in many paired comparison contexts, notable examples being the home-field effect in sport competitions or the white player advantage in chess. In paired-comparison models, such an order effect is typically described through the inclusion of an intercept term (e.g., Agresti, 2002),
$\Pr(Y_{ij}=1\mid\bm{\mu})=\Phi(\delta+\mu_{i}-\mu_{j}),$ (10)
@@ 3.2 Order effect and ties
Another common complication in paired comparisons is the presence of ties, represented by the null outcome $Y_{ij}=0$ whenever $i$ and $j$ are compared but neither beats the other. As discussed in Agresti (1992), ties can be incorporated in paired comparison models through a cumulative link model. The Thurstone-Mosteller model with an order effect corresponds to the latent continuous process $Z_{ij}=\delta+\mu_{i}-\mu_{j}+\epsilon_{ij},$ where $\epsilon_{ij}$ are uncorrelated standard normal variables. The outcome of the comparison depends on a threshold parameter $\gamma\geq 0$, so that $j$ beats $i$ if $Z_{ij}<-\gamma$, $i$ ties with $j$ if $-\gamma\leq Z_{ij}<\gamma$ and $i$ beats $j$ if $Z_{ij}\geq\gamma.$ The corresponding cumulative probabilities for the three outcomes given the strengths are
$\Pr(Y_{ij}\leq x\mid\bm{\mu})=\Phi(c_{x}-\mu_{i}+\mu_{j}),\quad x\in\{-1,0,1\},$
@@
As before, we proceed by empirical Bayes estimation of the model parameters $\gamma$, $\delta$ and $\lambda$ from the marginal distribution of the paired outcomes, and then compute the ridge estimates of the strength parameters. The marginal probabilities of the three possible outcomes are simply calculated by setting all strength parameters to zero, giving $\Pr(Y_{ij}=x)=\Phi(c_{x})-\Phi(c_{x-1})$, for $x\in\{-1,0,1\}$. From these marginal probabilities, consistent estimates of $\gamma$ and $\delta$ are readily obtained as
$\hat{\gamma}$ $=\frac{1}{2}\left\{\Phi^{-1}(1-\hat{p}_{1})-\Phi^{-1}(\hat{p}_{-1})\right\},$ (11)
$\hat{\delta}$ $=\frac{1}{2}\left\{\Phi^{-1}(\hat{p}_{1})-\Phi^{-1}(\hat{p}_{-1})\right\},$ (12)
where $\hat{p}_{-1}=(n+1)^{-1}\sum_{ij}1(Y_{ij}=-1)$ and $\hat{p}_{1}=(n+1)^{-1}\sum_{ij}1(Y_{ij}=1)$. In tournaments without ties, $\hat{\gamma}=0$ as expected.
Having estimated $\gamma$ and $\delta$, we proceed with estimation of the tuning parameter $\lambda$ by maximizing the pairwise log-likelihood (6), where $p_{rs}(\lambda)$ now denotes the joint probability of the outcomes of two paired comparisons with the same item appearing in either the first or the second position. For example, the probability that the first item $i$ obtains a win and a tie in such a couple of paired comparisons is
@@
The other bivariate probabilities involved in the pairwise likelihood are similarly computed. In these probabilities $\gamma$ and $\delta$ are replaced with their estimates (11)-(12) so that the pairwise likelihood depends on $\lambda$ only.
Maximization of the pairwise likelihood for $\lambda$ is conveniently performed by using the robust algorithm of Brent (2013, ch. 4) on the finite interval domain of $\tau$ from equation (7), with subsequent back-transformation as in equation (8). The same small-sample adjustment as before remains available through the pairwise likelihood penalty of equation (9).
@@ 4.1 True strengths normally distributed
The performance of the pairwise empirical Bayes method is illustrated first by simulations from the Thurstone-Mosteller model with normally distributed strength parameters. We considered 120 scenarios given by the combination of: three values for the number of items $p\in\{20,40,60\}$; eight values for the tuning parameter $\lambda\in\{2^{-1},2^{0},\ldots,\allowbreak 2^{6}\}$; and five sample sizes corresponding to five increasing fractions of the paired comparisons in a double round-robin tournament. More specifically, we simulated the full double-round-robin tournament and used the first $m$ rounds for training and the remaining $2(p-1)-m$ rounds for testing. The number of training rounds was chosen so as to approximate a specified proportion of the tournament paired comparisons. For example, when we consider 20% of paired comparisons for training and the number of items is $p=30$, the training set is taken to be the first 12 rounds because $0.2\times(2\times 29)=11.6$. The five sample fractions considered for training in our simulations are $20\%,30\%,40\%,50\%$ and $80\%$ of the paired comparisons in the double round-robin tournament. For each scenario, $1000$ replications were generated and analysed.
Figure 1: Logarithmic skill scores computed for incomplete double round-robin tournaments of different sizes. The columns correspond to various levels of incompleteness (20%, 30%, 50% or 80% of all possible paired comparisons used for training and the rest of paired comparisons for prediction). The rows correspond to different numbers of items (20, 40, 60). The item strengths are drawn from a zero-mean normal distribution with precision $\lambda$. The strengths are estimated by bias reduced maximum likelihood (BRMLE), maximum likelihood (MLE) and pairwise empirical Bayes (PEB). The reference method is a naive forecast that uses only the assumed marginal probability of a win for the first item and corresponds to a zero logarithmic skill score (grey solid line).
@@
The performance of our pairwise empirical Bayes method is compared here with three other methods: (1) maximum likelihood estimation, (2) bias-reduced maximum likelihood as in Firth (1993), and (3) maximum ridge-penalized likelihood with $\lambda$ determined by cross-validation. Bias-reduced maximum likelihood estimation is computed through the R (R Core Team, 2024) package brglm2 (Kosmidis, 2023). Since paired-comparison models are specified through pairwise differences of strengths, an arbitrary constraint is needed for identification of the maximum likelihood and bias-reduced maximum likelihood estimates; in the simulations we used the constraint $\mu_{1}=0$. Cross-validated ridge estimates were calculated via the glmnet package (Friedman et al., 2010; Tay et al., 2023), using tournament rounds as the assessment unit. Paraphrasing the typical language used in model validation, we could call such a design a ‘leave one round out’ cross-validation.
@@
We simulated paired comparisons without ties, in order to focus on the estimation of the strength parameters. Estimation of the tie threshold parameter $\gamma$ would not be problematic, but bias-reduced maximum likelihood and cross-validated logistic regression are computationally less expensive (for repeated use in a simulation experiment) in paired comparison models without ties. In our simulations the order effect parameter was set to $\delta=0.2$, meaning that in approximately 58% of paired comparisons the first item beats the second (since $\Phi(0.2)\approx 0.58$).
@@
The predictive performance of each method is measured by the classical logarithmic score (LS), defined as the negative average log-likelihood for the testing paired comparisons computed at the strengths estimated from the training comparisons. As reference we consider the naive forecast that predicts the result of every paired comparison using the assumed marginal probability of a win for the first item, that is $\text{LS}_{\text{naive}}=-0.58\log(0.58)-0.42\log(0.42)=0.68$. The naive forecast corresponds to predictions for the paired comparison model with strength parameters all zero, $\mu_{i}=0$ for all $i=1,\ldots,p$. With this reference, we define the logarithmic skill score as $\text{LSS}=1-\text{LS}/\text{LS}_{\text{naive}}=1-\text{LS}/0.68.$ The higher the logarithmic skill score, the better the predictive performance. Better forecasts on average than those obtained by the naive forecast have a logarithmic skill score above zero.
@@
We now comment on the results reported in Figure 1. As expected, as the number of items or the fraction of paired comparisons used for training increases, all methods converge to the same predictive performance. Furthermore, as $\lambda$ increases, estimates of item strengths converge to zero and therefore all methods produce predictions that are on average equivalent to those of the naive forecast. The predictive performance of maximum likelihood estimation is completely unsatisfactory when the training set is small and the number of items is 20, with its average logarithmic skill scores being worse than the naive forecast. The bias-reduced maximum likelihood estimator improves substantially on maximum likelihood in most settings, but with small training sets and small values of $\lambda$ it still leads to worse logarithmic skill scores on average than the naive forecast.
As shown in Figure 1, pairwise empirical Bayes has uniformly the best predictive performance of all methods considered. Pairwise empirical Bayes was even found to perform relatively well in a still more extreme setting, with $p=20$ items and only 4 paired comparisons per item (approximately 10% of the tournament) used for training; that extreme setting is not displayed in Figure 1 because maximum likelihood performs so poorly that its logarithmic skill score values fall too far below zero.